OpenStudy (anonymous):

Express this amount in dozens. (Use decimal form.) 7.5 cupcakes

OpenStudy (anonymous):

1 dozen = 12 ? dozen = 7.5 cross product you get 7.5*1 = 12*? so ? = 7.5/12 = 0.625 dozen

OpenStudy (anonymous):

thanks for helping!

OpenStudy (anonymous):

no problem!